OpenStudy (anonymous):

if 30 percent of 200 is equal to 250 percent of n, then n is equal to what number?

ganeshie8 (ganeshie8):

form an equation first

OpenStudy (anonymous):

(300/100)*100 = 250n / 100

OpenStudy (anonymous):

i mean ( 30/100)*100

ganeshie8 (ganeshie8):

30 percent of 200 = 250 percent of n

ganeshie8 (ganeshie8):

you mean (30/100) * 200 then its right. ..

OpenStudy (anonymous):

sooo 60 = 2.5n

OpenStudy (anonymous):

= 24

OpenStudy (anonymous):

\[\frac{3}{10}200=\frac{250}{100}n\]\[n=24 \]

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

thanksss youu

ganeshie8 (ganeshie8):

percent is just a ratio, as a fraction of 100 so we just replace percent by putting down 100